OpenStudy (anonymous):

PLZ HELP!!! Slove w^4-18w^2-2=0

OpenStudy (anonymous):

substitute w^2 as a

OpenStudy (anonymous):

?

OpenStudy (anonymous):

w^2=a then you will get quadratic equation

OpenStudy (anonymous):

w^4-18w^2-2=0 as he says substitute a for w^2 so a^2 = (w^2)^2 = w^4 equation reduces to a^2 - 18a - 2 = 0 this is a quadratic equation and you can find value of a then since a = w^2 put values of a = w^2 and take the root to get value of w

OpenStudy (anonymous):

So the answer is?

OpenStudy (anonymous):

42